For a real gas R=0.08 and T=300 . d/P=4+0.04P+0.002P2+0.0001P3....... 
Find molar mass of the gas.

Grade:

1 Answers

Vikas TU
5 years ago
Dear Student,
For P very very small ,
d/P = 4.
And density .i.e. d = PM/RT
Or
d/P = M/RT  = 4 
M = 4RT => 4*0.08*300 => 69 g